OpenStudy (anonymous):

Please use the derivative to show why f(x)=integral 1/sqrt(4+e^t)dt from 0 to lnx, with x>0, has an inverse function

OpenStudy (anonymous):

did you find the derivative?

OpenStudy (anonymous):

maybe i should ask "do you know how to find the derivative?"

OpenStudy (anonymous):

no, sorry, I haven't found the derivative.

OpenStudy (anonymous):

I actually don't know how to find the derivative.

OpenStudy (anonymous):

the derivative of the integral is the integrand that, and the chain rule

OpenStudy (anonymous):

\[\frac{d}{dx}\int_a^x f(t)dt=f(x)\]

OpenStudy (anonymous):

\[\frac{d}{dx}\int_a^{g(x)}f(t)dt=f(g(x))g'(x)\]

OpenStudy (anonymous):

you have \[\int_0^{\ln(x)}\frac{1}{\sqrt{4+e^t}}dt\] if i am reading it correctly

OpenStudy (anonymous):

replace \(t\) by \(\ln(x)\) in \(\frac{1}{\sqrt{4+e^t}}\) and multiply by \(\frac{1}{x}\) to get the derivative

OpenStudy (anonymous):

that makes your derivative \[\large \frac{1}{\sqrt{4+e^{\ln(x)}}}\times \frac{1}{x}\] or more simply \[\frac{1}{x\sqrt{4+x}}\]

OpenStudy (anonymous):

Thank you so much! that makes a lot of sense. I actually understand how to do it now!